Name the complex conjugate, then find product of the complex number with its conjugate.
3-i
answers:
A) -3+i;10
B) 3+i;8
C) 3+i;10
D) -3+i;8
: Name the complex conjugate, then find product of the complex number with its conjugate.
3-i
answers:
A) -3+i;10
B) 3+i;8
C) 3+i;10
D) -3+i;8

Answer by Earlsdon(3748)

The complex conjugate of a complex number, (a+bi) is just (a-bi), so for:
(3-i) the complex conjugate is (3+i) and the product of these two can be found using the FOIL method:
(3-i)(3+i) = 9+3i-3i-i^2 Simplify the product:
9-i^2 Substitute i^2 = -1
9-(-1) = 9+1 = 10
Answer C: (3+i), 10